Find the equation of the ellipse whose focus is (1,0), the directrix is x+y+1=0 and eccentricity is 1/(square root)2.

Grade:11

1 Answers

Nishant Vora IIT Patna
7 years ago
Use the basic definition,

PS = e PM

PS^2 = e^2 PM^2

(h-1)^2 + k^2 = \frac{1}{2} (\frac{h+k+1}{\sqrt{2}})^2